Rank the following in order of weakest to strongest intermolecular forces:
Answer #1

Answer:

Here first (Left) molecule represents the Ethoxymethane = CH3-CH2-O-CH3 (Ether)

Second (middle) Acetone = CH3-CO-CH3 (Ketone)

Third (Right) Iso-propyl alcohol = CH3-CH(OH)-CH3 (Alcohol)

Explanation:

Here the strongest inter molecular forces are present in the alcohol due to the presence of "-O-H---" hydrogen bonding (due to the presence of O-H bond).

Now we are left with Ketone and Ether functional groups. Here, both the functional groups results in dipole-dipole interactions (due to difference in electronegativity between the C-O bond). But inter molecular forces are stronger in Ketones.

Reason: The double bond (ketone) between C=O is more polar compare to the single bond present in C-O (ether).

Order of intermolecular forces: Third (alcohol) > Second (ketone) > First (ether)

Please let me know, if you need further help for the same by commenting below the answer.

Thanks